let f(X)=5x^3-13x^2+11x+4 if the interval of difference in 1 then cosider following statement 1. \Delta ^3f(x)=30 for all x 2. \Delta ^2f(x)=34 for x=1 3 .\Deltaf(x)= 4 foe x=1 which of the statement  c. if \Delta ^nO^m = \Delta ^nX^m then what is the value of O^6. what is the fun where first difference in 3x^2+7x+10 and assume h=1.
Answer #1

above statemer. let f@= 5x3-132²H1X+4. if the interval of difference in Lither compides the following statement, ® 13 f(1) =Caic f(at2) 5(472)3 - 13(2+2)? +1 (442) ty 5013 + 12 +1971 tiy no seati) = 5fantoj - 13/4+1) + 11(21+0+y sa²+ 20²7 and son 57let the required fun. 121 € (2) given: AfC00 3262+75 +10 let 32827721710 = 32P+AS+B = 326421-1+A1217 B TB 2130 B 10 2-1 A + Bo - ND= E-1 given som = 3m loco then 36 30/000 (E-1} lase Ef(x) = futh)- £²5(1) = f(a+gh) 3E73E-1270 (x+3h) – 3(3+297 +3(0274